OpenStudy (anonymous):

Solve the equation 2 cos X - sec x = 1 on the interval [0,2pi)

OpenStudy (anonymous):

Step 1: Change sec x = 1/cos x (as sec x and cos x are receiprocals of each other).

OpenStudy (anonymous):

Tell me when you did that.

OpenStudy (anonymous):

done

OpenStudy (anonymous):

Step 2: Multiply both sides of the equation by cos x. (This will get rid of the fraction's denominator on the left side). Tell me when you have done that.

OpenStudy (anonymous):

done

OpenStudy (anonymous):

what do you have as your equation now?

OpenStudy (anonymous):

2 cos^2 x - 1 = cos x

OpenStudy (anonymous):

You should have: 2 cos^2 (x) - 1 = cos x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Subtract cos x from both sides...we have a quadratic equation... so you have 2 cos^2(x) - cos x - 1 = 0

OpenStudy (anonymous):

We will now factor that quadratic equation: (2 cos x + 1)(cos x - 1) = 0

OpenStudy (anonymous):

Tell me when you clearly understand everything till this point.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We now set each factor equal to 0: 2 cos x + 1 = 0 2 cos x = -1 cos x = -1/2 cos x - 1 = 0 cos x = 1

OpenStudy (anonymous):

now you dind the angles.

OpenStudy (anonymous):

-pi/6 only?

OpenStudy (anonymous):

cos x = 1 at x = 0

OpenStudy (anonymous):

cos x = -1/2 at 2pi/3 and 4pi/3

OpenStudy (anonymous):

3 solutions